A JFET has IDSS= 20 mA. At what gate source voltage (VGS) will it`s drain current be 5 mA, it`s pinch off voltage is 4.2 V

Grade:12th pass

1 Answers

Khimraj
6 years ago
Formula of drain current
ID = IDSS(1 – VGS/VP)2
Given that ID = 5mA, IDSS = 20mA, VP= 4.2 V
ID/IDSS = (1 – VGS/4.2)2
¼ = (1 – VGS/4.2)2
½ = (1 – VGS/4.2)
0.5*4.2 = 4.2 – VGS
VGS= 4.2 – 2.1 = 2.1V.
Hope it clears. If you like answer then please approve it.